Joe is 6 years older than Sam In 5 years time their combined ages will be 36 how old are they now
Answer by ankor@dixie-net.com(22740)

Write an equation for each statement
:
Joe is 6 years older than Sam
j = s+6
j - s = 6
In 5 years time their combined ages will be 36
(j+5) + (s+5) = 36
j + s + 10 = 36
j + s = 36 - 10
j + s = 26
Combine with the 1st equation, use elimination
j - s = 6
j + s = 26
-------------Adding eliminates s, find j
2j = 32
j = 16 is Joe's age
:
how old are they n)w
I'll let you find Sam's age, Check your solutions in the 2nd statement